OpenStudy (anonymous):

Solve for x. 5x - 11 = 7x - 3

OpenStudy (anonymous):

first collect like terms

OpenStudy (anonymous):

which is move x to other side and then numbers to the other side

OpenStudy (anonymous):

x=-4 Also when ever you need answers this website is VERY helpful http://www.wolframalpha.com/

OpenStudy (nurali):

5x - 11 = 7x - 3 5x-7x=-3+11 -2x=8 x=-8/2 x=-4